without graphing, how to state the slope and the y-
intercept of each line whose equation is given?

Answer by mananth(16946)

Assume the equation is 3x+y=6
convert to y= mx+b form where m is the slope & b is the y intercept
y=-3x+6
compare with y = mx+b
m=-3 & b=6
slope = 3 & y intercept = 6